How much Cuso4 . 5 H2o is required to prepare 500 c.c if 0.5 N solution

Grade:12

2 Answers

varun
6 years ago
HELLO PRINCE,
Molecular weight of CuSO4.5H2O 
=> 63.5 + 32 + 4X16 + 5X18
=>249.5
To prepare 1N solution
we have to dissolve ,= 249/5/2 gm
                                = 124.75 gm
so to prepare 500 cc ,1N solution, 124.75/2
=>62.375 gm
So to prepare 500 cc 0.5N soluiton
                                                        = 62.375 X 0.5
                                                         =  31.18 gm is required 
this is the simple and shortcut method 
THANK YOU
 
Patim y
3 years ago
We have 
Equivalent weight= 249.5/2 =124.75
Now using
W=VEN/1000 
We get 
W= 500 × 0.5×2×124.75/ 1000 
W= 31.18 gm